OpenStudy (lemanual):

Help please...Using the completing-the-square method, find the vertex of the function f(x) = 2x2 − 8x + 6 and indicate whether it is a minimum or a maximum and at what point. (2 points) Maximum at (2, −2) Minimum at (2, −2) Maximum at (2, 6) Minimum at (2, 6)

OpenStudy (misty1212):

HI!!

OpenStudy (lemanual):

Hello : )

OpenStudy (misty1212):

it is a parabola that opens up, so the second coordinate of the vertex is a minimum (lowest point on the curve)

OpenStudy (misty1212):

so it is either B or D to choose, plug in 2 for x and see if you get -2 or 6

OpenStudy (lemanual):

Ok ty

OpenStudy (lemanual):

Can I ask 2 more?

OpenStudy (lemanual):

@misty1212 Can I ask 2 more please?

OpenStudy (misty1212):

sure

OpenStudy (lemanual):

Rewrite f(x) = 3(x − 2)2 + 1 from vertex form to standard form. (2 points) f(x) = 3x2 + 13 f(x) = 3x2−12x + 13 f(x) = 3x2 + 12x − 11 f(x) = 9x2− 36x + 37

OpenStudy (lemanual):

(For this one I'm checking to see if this is right) Using the completing-the-square method, rewrite f(x) = x2 + 4x − 1 in vertex form. (2 points) f(x) = (x + 2)2 + 1 f(x) = (x + 2)2 f(x) = (x + 2)2 + 4 f(x) = (x + 2)2−5 <--- My answer

OpenStudy (misty1212):

second one is right for sure

OpenStudy (lemanual):

Ok

OpenStudy (misty1212):

\[ f(x) = 3(x − 2)^2 + 1=2(x-2)(x-2)+1=3(x^2-4x+4)-1\] etc

OpenStudy (misty1212):

in other words, expand and combine like terms

OpenStudy (misty1212):

i think you get \[3x^2-12x+11\] if i did it right

OpenStudy (misty1212):

ooopps n

OpenStudy (misty1212):

\[3x^2-12x+13\] looks better

OpenStudy (lemanual):

Ty so much Sorry Took me so long to answer back